# The anomalous magnetic moment of the muon: status and perspectives

### v1

@@ 1 INTRODUCTION
[…]
where $c_{L,R}^{\ell}$ are ${\mathcal{O}}(1)$ couplings (with the superscript indicating that the BSM physics need not be lepton flavor universal), $\Lambda_{\text{BSM}}$ refers to the BSM scale, the factor $16\pi^{2}$ is generic for a quantum-loop correction, and $R_{\chi}$ accounts for the nature of the chirality flip. The latter can induce sizable so-called chiral enhancements [1, 2], e.g., when the chirality flip is mediated by a new heavy fermion $F$, $R_{\chi}=m_{F}/m_{\ell}$, while another special case, that of radiative mass generation, would amount to $c_{L}^{\ell}c_{R}^{\ell}R_{\chi}\simeq 16\pi^{2}\simeq 160$. Further scenarios yielding contributions well above the benchmarks defined in Eq. (1) can be obtained in the presence of light new particles [1].
In the case of the electron, the comparison of direct measurement [3] and SM prediction [4, 5, 6, 7, 8, 9, 10, 11] constrains BSM effects to $|a_{e}^{\text{BSM}}|\lesssim 1.0\times 10^{-12}$. However, at present the sensitivity is limited by a $5.5\sigma$ tension in the fine-structure constant $\alpha$ [12, 13], whose resolution would propel the reach to $1.5\times 10^{-13}$. The next generation of atom-interferometry experiments currently ongoing should be able to achieve this goal, and together with an improved direct measurement, which is also in progress, could improve the sensitivity by another factor of $10$, only an order of magnitude away even from the quadratic benchmark (1).
@@
dominated by the FNAL E989 experiment, already lies within a factor of two of the benchmark in Eq. (1), and thus would allow one to exclude wide classes of BSM scenarios and parameter space corresponding to the myriad enhancements beyond the naive quadratic scaling at $\Lambda_{\text{BSM}}=1\,\text{TeV}$. Unfortunately, this sensitivity is currently significantly diluted by a larger uncertainty in the SM prediction [22]
@@
Figure 1: A: Experimental results from BNL (blue triangles) [75, 76, 77], [^2] The quoted BNL ${\mathcal{R}}_{\mu}^{{}^{\prime}}$ and $a_{\mu}$ values were computed by the FNAL Muon g2g-2 collaboration [14], converting the reported values for free protons in vacuum in the BNL references and using the converted ${\mathcal{R}}_{\mu}^{{}^{\prime}}$ values to recompute the $a_{\mu}$ values relying on CODATA 2022 [79] for the external inputs.FNAL (red squares) [17, 15, 14], and the world average (black diamond). The uncertainties combine statistics and systematics. BNL values are rounded to the nearest $10\,\text{ppb}$. All measurements employed positive muons except BNL-01, which ran with negative muons (inverted triangle). The CERN III result [78], with a precision of $7\,300\,\text{ppb}$, is not shown. B: Final relatively balanced systematic uncertainties at FNAL, see Sec. 3 for a detailed discussion of the individual contributions.
In this review, we focus on the experimental techniques and innovations that led to Eq. (2), reflecting the progression from the Brookhaven National Laboratory (BNL) to the FNAL experiment as summarized in Fig. 2(A), see Secs. 2 and 3. We then summarize the current status of the SM prediction leading to Eq. (3), see Sec. (4). In Sec. 5, we describe ongoing efforts to improve the SM prediction to try and match the experimental precision, and contemplate future experiments that could surpass the current $124\,\text{ppb}$.
@@ 2.1.1 Miracle #1: The intrinsic properties of the muon
where $E_{r}$ is the positron energy and $\theta$ is the angle between its momentum and the muon spin. The asymmetry function $A_{r}(E)$ represents the strength and sign of the correlation; the key will be that the highest-energy positrons are preferentially emitted in the direction of the positive muon’s spin. Both predictions were promptly confirmed [83], demonstrating a violation of parity in the weak interaction and establishing the fundamental principles needed to measure the magnetic moment.
@@ 2.1.3 Miracle #3: The magic $\gamma$
To reach high precision, muons must be stored in the ring for approximately 10 lifetimes. A system of electrostatic quadrupole (ESQ) plates placed inside the SR provides weak vertical focusing. The SR thus resembles a large Penning trap; see Fig. 2(E). The measured precession frequency $\omega_{a}^{m}$ will differ from the value needed to compute $a_{\mu}$ because of the tiny influence of nonplanar trajectories and the presence of the electric field $\mathbf{E}$}. A more complete [^5] An additional perturbation to $\omega_{a}$ is required for a nonzero electric dipole moment, which has the form $\frac{\eta}{2}({\boldsymbol{\beta}}\times{\mathbf{B}}+{\mathbf{E}})$. expression for $\omega_{a}$ that includes these effects is given by [84]
@@ 3 THE FERMILAB EXPERIMENT
The FNAL-E989 experiment operated for six data-taking periods (Runs 1–6), all using positive muons. The main elements of the BNL equipment reused at FNAL include the magnetic SR [87], the superconducting inflector magnet [88], the vacuum chamber system, and the ESQ plates. The fast kicker and ESQ power supplies are new, as are all detectors, electronic systems, precession and field data acquisition systems, and the pulsed-NMR probes and much of the electronics. The g2g-2 results are published in three Physical Review Letters [17, 15, 14]. Detailed reports on the methods to determine the precession frequency, beam dynamics, and magnetic field can be found in the Physical Review [18, 20, 19, 16].
@@ 3.1 Innovative aspects of the FNAL E989 experiment
Figure 2: A: Muon Campus beamline overview and B: schematic of key elements in the SR. C: The beam intensity (orange) and kicker strength (blue) profiles vs. time at injection. The resulting storage efficiency, plotted with respect to mean momentum vs. time is shown in the lower panel. D: The intensity of events striking a calorimeter at early (blue) and later (red) times. The peaks align with $\omega_{c}$ but the width evolves owing to the finite momentum spread in the bunch. Note, $\omega_{a}$ and the exponential decay are divided out in this figure. E: Momentum distribution (upper axis) of the stored beam overlaid with the ESQ plate positions and their electric field lines. The collimator is shown as a black ring.
@@ 3.1.1 Injecting and storing polarized muons
The muon beam enters the SR through a hole in the back-leg of the magnet yoke, then passes through a 1.7 m-long narrow superconducting inflector magnet [88] that creates a field-free corridor into the storage region. Figure 2(B) illustrates the locations of key components related to muon storage and the $\omega_{a}$ measurement. A newly developed pulsed magnetic kicker system (K1, K2, K3) provides the needed angular deflection [89]. Figure 2(C) illustrates the challenge of this process. The vertical dashed gray lines are separated by the 149 ns cyclotron period. The kicker pulse strength (blue) is overlaid with a representative muon bunch intensity profile (orange). The provided deflection is not uniform. Although the average of the pulse train can be nearly centered at the magic radius, $x=0$ in the figure, the storage efficiency and average momentum that is stored varies from the beginning to the end of the bunch train. The lower panel illustrates the realized variation in intensity and central momentum vs. time along the bunch train. Nevertheless, approximately 4000 muons are stored each “fill” with a momentum width $\delta P/P_{0}\simeq 0.15\%$, and a mean momentum nearly centered on the magic momentum.
@@ 3.1.2 Measuring the precession frequency
The positrons having higher energies form a spectrum $N(t)$ as shown in Fig. 3(A). A fit to $N(t)$ determines $\omega_{a}^{m}$. Events entering the spectrum are weighted by an effective, measured energy-dependent asymmetry, $A(E)$, which varies smoothly from $0.0$ to $0.8$ across the 1.0 to 3.1 GeV interval of included positron energies. An innovative system [93] of pulsed lasers flashes each crystal in patterns to establish long- and short-term gain corrections functions that are applied during reconstruction of the raw data. The stability of the energy response is therefore maintained throughout a fill to a few parts in $10^{-4}$ to avoid time dependence in $\phi(E)$.
@@ 4 STATUS OF THE STANDARD-MODEL PREDICTION
Figure 5: A: Summary of various determinations of $a_{\mu}^{\text{HVP, LO}}$, propagated to $a_{\mu}^{\text{SM}}$. The first two panels refer to data-driven determinations, where the three points for each $e^{+}e^{-}$ experiment reflect the CHKLS [99, 100, 101, 102] (triangle), DHMZ [103, 104, 105] (diamond), and KNTW [106, 8, 107] (square) methods, see WP25 for details. The gray band indicates the WP20 result, based on the $e^{+}e^{-}$ experiments above the first dashed line. The second panel represents $e^{+}e^{-}$ experiments that became available afterwards as well as the $\tau$-based estimate from WP25. The last panel summarizes lattice-QCD determinations, including the hybrid evaluation by BMW/DMZ-24 [38], three individual lattice-QCD calculations by RBC/UKQCD-24+18 [40], Mainz/CLS-24 [41], BMW-20 [30], and five lattice HVP averages from WP25. The blue band refers to the final WP25 result, which coincides with “Avg. 1.” In all cases, except for the gray WP20 band, the remaining contributions to $a_{\mu}^{\text{SM}}$ beyond $a_{\mu}^{\text{HVP, LO}}$ are taken from WP25. The red band denotes the experimental world average. Adapted from Ref. [22]. B: Sample diagrams for $a_{\mu}$ in the SM: solid lines refer to the muon, wiggly lines to photons, and the red blobs to hadronic matrix elements. The first diagram gives the leading QED contribution by Schwinger, the second one represents a one-loop EW diagram with $Z$ exchange, and the last two diagrams correspond to HVP and HLbL topologies, respectively. C: Summary of HLbL evaluations, from data-driven methods (green), lattice QCD (blue), and combinations (black). The averages are from WP20 and WP25, respectively, the other points refer to HSZ-24 [108, 66], RBC/UKQCD-19 [70], Mainz/CLS-21+22 [71, 72], RBC/UKQCD-23 [73], and BMW-24 [74]. Adapted from Ref. [22].
@@
where the latter materialize from quark loops, which are, in principle, described by quantum chromodynamics (QCD), but at low energies require nonperturbative techniques because the quarks and gluons are confined into hadrons. They are further separated into topologies featuring hadronic vacuum polarization (HVP) and hadronic light-by-light scattering (HLbL), see Fig. 5(B) for representative diagrams in each category. Predicting $a_{\mu}^{\text{SM}}$ at the level ${\mathcal{O}}(10^{-10})$ requires exquisite control over all parts of the SM, and it is quite remarkable that this is indeed possible. In the following we briefly discuss the status for each class of corrections, referring to Refs. [109, 22] for a thorough review, labeled as WP20 [109] and WP25 [22] below. [^9] These White Papers (WP) were compiled by the Muon g2g-2 Theory Initiative [110], reflecting a community consensus of the status of the SM prediction in 2020 and 2025, respectively. The resulting bottom line is summarized in Table 2.
Table 2: Summary of all contributions to the SM prediction and comparison to experiment, adapted from Ref. [22].
@@ 4.3 Hadronic vacuum polarization
and a kernel function $\hat{K}(s)$, establishing a direct link to experimental data for hadronic cross sections. In particular, the leading-order (LO) contribution is defined photon-inclusively, but removing vacuum-polarization effects in the photon propagator. Accordingly, the integration starts at the threshold of the $e^{+}e^{-}\to\pi^{0}\gamma$ channel, $s_{\text{thr}}=M_{\pi^{0}}^{2}$. In WP20, the final result for the HVP contribution was based on $e^{+}e^{-}$ data, including data for the by far numerically dominant $e^{+}e^{-}\to\pi^{+}\pi^{-}$ channel from the SND06 [140], CMD-2 [141], BaBar [142], KLOE [143], and BESIII [144] experiments. Moreover, different methodologies for the compilation of $e^{+}e^{-}$ data were taken into account [103, 106, 99, 145, 104, 8], and an additional uncertainty beyond a simple scale factor was introduced to account for the tension between the BaBar and KLOE measurements, leading to the gray band in Fig. 5(A). Subsequently, new $\pi^{+}\pi^{-}$ measurements from SND20 [146] and CMD-3 [147, 148] have become available, and while the former falls within the previous range, the latter increases tensions to a level that cannot be taken into account anymore by a meaningful inflation of uncertainties. Due to these severe tensions, combined with the fact that most experiments do not cover the entire phase space, already different methodologies to assign HVP integrals for a given $\pi^{+}\pi^{-}$ experiment can lead to different results, as illustrated by the triangle, diamond, and square points in Fig. 5(A).

### v2

@@ 1 INTRODUCTION
$a_{\ell}^{\text{BSM}}\simeq R_{\chi}\frac{c_{L}^{\ell}c_{R}^{\ell}}{16\pi^{2}}\times\frac{m_{\ell}^{2}}{\Lambda_{\text{BSM}}^{2}}\simeq\big(c_{L}^{\ell}c_{R}^{\ell}R_{\chi}\big)\bigg(\frac{1\,\text{TeV}}{\Lambda_{\text{BSM}}}\bigg)^{2}\times\begin{cases}1.7\times 10^{-15}\qquad&\ell=e\,,\\ 7.1\times 10^{-11}\qquad&\ell=\mu\,,\\ 2.0\times 10^{-8}\qquad&\ell=\tau\,.\end{cases}$ (1)
Setting the prefactor on the right-hand side equal to $1$ provides a useful benchmark for evaluating the power of an experimental constraint. In Eq. (1) the $c_{L,R}^{\ell}$ are ${\mathcal{O}}(1)$ couplings (with the superscript indicating that the BSM physics need not be lepton flavor universal), $\Lambda_{\text{BSM}}$ refers to the BSM scale, the factor $16\pi^{2}$ is generic for a quantum-loop correction, the $m_{\ell}^{2}$ dependence accounts for the chirality flip required in the magnetic moment operator, and $R_{\chi}$ is a factor to change the value if the chirality flip has a different origin. In specific models, $R_{\chi}$ can induce so-called chiral enhancements [1, 2], e.g., when the chirality flip is mediated by a new heavy fermion $F$, $R_{\chi}=m_{F}/m_{\ell}$, while another special case, that of radiative mass generation, would amount to $c_{L}^{\ell}c_{R}^{\ell}R_{\chi}\simeq 16\pi^{2}\simeq 160$. Further scenarios yielding contributions well above the benchmark (1) can be obtained in the presence of light new particles [1].
In the case of the electron, the comparison of direct measurement [3] and SM prediction [4, 5, 6, 7, 8, 9, 10, 11] constrains BSM effects to $|a_{e}^{\text{BSM}}|\lesssim 1.0\times 10^{-12}$. However, at present the sensitivity is limited by a $5.5\sigma$ tension in the fine-structure constant $\alpha$ [12, 13], whose resolution would propel the reach to $1.5\times 10^{-13}$. The next generation of atom-interferometry experiments currently ongoing should be able to achieve this goal, and together with an improved direct measurement, which is also in progress, could improve the sensitivity by another factor of $10$, only an order of magnitude away even from the benchmark (1).
@@
dominated by the FNAL E989 experiment, already lies within a factor of two of the benchmark (1), and thus would allow one to exclude wide classes of BSM scenarios and parameter space corresponding to the myriad enhancements beyond the naive quadratic scaling at $\Lambda_{\text{BSM}}=1\,\text{TeV}$. Unfortunately, this sensitivity is currently significantly diluted by a larger uncertainty in the SM prediction [22]
@@
Figure 1: A: Experimental results from BNL (blue triangles) [75, 76, 77], [^2] The quoted BNL $a_{\mu}$ values were computed by the FNAL Muon g2g-2 collaboration [14], converting the published BNL-E821 results using updated values from CODATA 2022 [79] for the external inputs of fundamental constants.FNAL (red squares) [17, 15, 14], and the world average (black diamond). The uncertainties combine statistics and systematics. BNL values are rounded to the nearest $10\,\text{ppb}$. All measurements employed positive muons except BNL-01, which ran with negative muons (inverted triangle). The CERN III result [78], with a precision of $7\,300\,\text{ppb}$, is not shown. B: Final relatively balanced systematic uncertainties at FNAL, see Sec. 3 for a detailed discussion of the individual contributions.
In this review, we focus on the experimental techniques and innovations that led to Eq. (2). In particular, we focus on the evolution from the CERN g2g-2 experiment CERN-III to the Brookhaven National Laboratory (BNL) experiment BNL-E821 and then to the FNAL experiment FNAL-E989, as summarized in Fig. 2(A), see Secs. 2 and 3. We then summarize the current status of the SM prediction leading to Eq. (3), see Sec. (4). In Sec. 5, we describe ongoing efforts to improve the SM prediction to try and match the experimental precision, and contemplate future experiments that could surpass the current $124\,\text{ppb}$.
@@ 2.1.1 Miracle #1: The intrinsic properties of the muon
where $E_{r}$ is the positron energy and $\theta$ is the angle between its momentum and the muon spin. The asymmetry function $A_{r}(E)$ represents the strength and sign of the correlation; the key will be that the highest-energy positrons are preferentially emitted in the direction of the positive muon’s spin. Both predictions were promptly confirmed [83, 84], demonstrating a violation of parity in the weak interaction and establishing the fundamental principles needed to measure the magnetic moment.
@@ 2.1.3 Miracle #3: The magic $\gamma$
To reach high precision, muons must be stored in the ring for approximately 10 lifetimes. A system of electrostatic quadrupole (ESQ) plates placed inside the SR provides weak vertical focusing. The SR thus resembles a large Penning trap; see Fig. 2(E). We use $\omega_{a}^{m}$ to denote the precession frequency actually measured in the experiment. This differs from the nominal precession frequency because of the tiny influence of nonplanar trajectories and the presence of the electric field $\mathbf{E}$}. A more complete [^5] An additional perturbation to $\omega_{a}$ is required for a nonzero electric dipole moment, which has the form $\frac{\eta}{2}({\boldsymbol{\beta}}\times{\mathbf{B}}+{\mathbf{E}})$. expression for $\omega_{a}$ that includes these effects is given by [85]
@@ 3 THE FERMILAB EXPERIMENT
The FNAL-E989 experiment operated for six data-taking periods (Runs 1–6), all using positive muons. The main elements of the BNL equipment reused at FNAL include the magnetic storage ring (SR) [88], the superconducting inflector magnet [89], the vacuum chamber system, and the electrostatic quadrupole (ESQ) plates. The fast kicker and ESQ power supplies are new, as are all detectors, electronic systems, precession and field data acquisition systems, and the pulsed-NMR probes and much of the electronics. The g2g-2 results are published in three Physical Review Letters [17, 15, 14]. Detailed reports on the methods to determine the precession frequency, beam dynamics, and magnetic field can be found in the Physical Review [18, 20, 19, 16].
@@ 3.1 Innovative aspects of the FNAL E989 experiment
Figure 2: A: Muon Campus beamline overview and B: schematic of key elements in the SR. C: The entering beam intensity (brown) and kicker strength (blue) profiles vs. time of injection. The stored beam profile (solid orange) is the product of the $t_{0}$ profile and the overall storage efficiency (black, bottom) calculated from special systematic runs. D: The intensity of events striking a calorimeter at early (blue) and later (red) times. The peaks align with $\omega_{c}$ but the width evolves owing to the finite momentum spread in the bunch. Note, $\omega_{a}$ and the exponential decay are divided out in this figure. E: Momentum distribution (upper axis) of the stored beam overlaid with the ESQ plate positions and their electric field lines. The collimator is shown as a black ring.
@@ 3.1.1 Injecting and storing polarized muons
The muon beam enters the SR through a hole in the back-leg of the magnet yoke, then passes through a 1.7 m-long narrow superconducting inflector magnet [89] that creates a field-free corridor into the storage region. Figure 2(B) illustrates the locations of key components related to muon storage and the $\omega_{a}$ measurement. A newly developed pulsed magnetic kicker system (K1, K2, K3) provides the needed angular deflection [90]. Figure 2(C) illustrates the challenge of this process. The vertical dashed lines are separated by the 149 ns cyclotron period. The kicker pulse strength (blue) is overlaid with a representative muon bunch intensity profile (brown). The resulting stored profile (orange) is not uniform and differs slightly bunch to bunch. The bottom panel illustrates the relative efficiency vs. time for this example. Although the average of the pulse train can be nearly centered at the magic radius, $x=0$ in the figure, the relative storage efficiency varies from the beginning to the end of the bunch train. Nevertheless, approximately 4000 muons are stored each “fill” with a momentum width $\delta P/P_{0}\simeq 0.15\%$, and a mean momentum nearly centered on the magic momentum.
@@ 3.1.2 Measuring the precession frequency
The positrons having higher energies form a spectrum $N(t)$ as shown in Fig. 3(A). A fit to $N(t)$ determines $\omega_{a}^{m}$, Eq. (7). Events entering the spectrum are weighted by an effective, measured energy-dependent asymmetry, $A(E)$, which varies smoothly from $0.0$ to $0.8$ across the 1.0 to 3.1 GeV interval of included positron energies. An innovative system [94] of pulsed lasers flashes each crystal in patterns to establish long- and short-term gain corrections functions that are applied during reconstruction of the raw data. The stability of the energy response is therefore maintained throughout a fill to a few parts in $10^{-4}$ to avoid time dependence in $\phi(E)$.
@@ 4 STATUS OF THE STANDARD-MODEL PREDICTION
Figure 5: A: Summary of various determinations of $a_{\mu}^{\text{HVP, LO}}$, propagated to $a_{\mu}^{\text{SM}}$. The first two panels refer to data-driven determinations, where the three points for each $e^{+}e^{-}$ experiment reflect the CHKLS [100, 101, 102, 103] (triangle), DHMZ [104, 105, 106] (diamond), and KNTW [107, 8, 108] (square) methods, see WP25 for details. The gray band indicates the WP20 result, based on the $e^{+}e^{-}$ experiments above the first dashed line. The second panel represents $e^{+}e^{-}$ experiments that became available afterwards as well as the $\tau$-based estimate from WP25. The last panel summarizes lattice-QCD determinations, including the hybrid evaluation by BMW/DMZ-24 [38], three individual lattice-QCD calculations by RBC/UKQCD-24+18 [40], Mainz/CLS-24 [41], BMW-20 [30], and five lattice HVP averages from WP25. The blue band refers to the final WP25 result, which coincides with “Avg. 1.” In all cases, except for the gray WP20 band, the remaining contributions to $a_{\mu}^{\text{SM}}$ beyond $a_{\mu}^{\text{HVP, LO}}$ are taken from WP25. The red band denotes the experimental world average. Adapted from WP25. B: Sample diagrams for $a_{\mu}$ in the SM: solid lines refer to the muon, wiggly lines to photons, and the red blobs to hadronic matrix elements. The first diagram gives the leading QED contribution by Schwinger, the second one represents a one-loop EW diagram with $Z$ exchange, and the last two diagrams correspond to HVP and HLbL topologies, respectively. C: Summary of HLbL evaluations, from data-driven methods (green), lattice QCD (blue), and combinations (black). The averages are from WP20 and WP25, respectively, the other points refer to HSZ-24 [109, 66], RBC/UKQCD-19 [70], Mainz/CLS-21+22 [71, 72], RBC/UKQCD-23 [73], and BMW-24 [74]. Adapted from WP25.
@@
where the latter materialize from quark loops, which are, in principle, described by quantum chromodynamics (QCD), but at low energies require nonperturbative techniques because the quarks and gluons are confined into hadrons. They are further separated into topologies featuring hadronic vacuum polarization (HVP) and hadronic light-by-light scattering (HLbL), see Fig. 5(B) for representative diagrams in each category. Predicting $a_{\mu}^{\text{SM}}$ at the level ${\mathcal{O}}(10^{-10})$ requires exquisite control over all parts of the SM, and it is quite remarkable that this is indeed possible. In the following, we discuss the status for each class of corrections. Our discussion is strongly based on the White Papers (WP) of the Muon g2g-2 Theory Initiative [110], to which we will refer as WP20 [111] and WP25 [22] throughout this section, each reflecting a community consensus at the respective time. The resulting bottom line is summarized in Table 2.
Table 2: Summary of all contributions to the SM prediction and comparison to experiment, adapted from WP25.
@@ 4.3 Hadronic vacuum polarization
and a kernel function $\hat{K}(s)$, establishing a direct link to experimental data for hadronic cross sections. At the required level of precision, it is important to be precise about the QED effects included in this quantity. $R_{\text{had}}(s)$ does not include QED corrections to the photon propagator, but it does include final states with photons in addition to hadrons. It also includes effects of virtual photons coupling to quarks, but not virtual photons also coupling to the initial state $e^{+}e^{-}$. Accordingly, the integration in Eq. (20) starts at the threshold of the $e^{+}e^{-}\to\pi^{0}\gamma$ channel, $s_{\text{thr}}=M_{\pi^{0}}^{2}$. In WP20, the final result for the HVP contribution was based on $e^{+}e^{-}$ data, including data for the by far numerically dominant $e^{+}e^{-}\to\pi^{+}\pi^{-}$ channel from the SND06 [141], CMD-2 [142], BaBar [143], KLOE [144], and BESIII [145] experiments. Moreover, different methodologies for the compilation of $e^{+}e^{-}$ data were taken into account [104, 107, 100, 146, 105, 8], and an additional uncertainty beyond a simple scale factor was introduced to account for the tension between the BaBar and KLOE measurements, leading to the gray band in Fig. 5(A). Subsequently, new $\pi^{+}\pi^{-}$ measurements from SND20 [147] and CMD-3 [148, 149] have become available, and while the former falls within the previous range, the latter increases tensions to a level that cannot be taken into account anymore by a meaningful inflation of uncertainties. Due to these severe tensions, combined with the fact that most experiments do not cover the entire phase space, already different methodologies to assign HVP integrals for a given $\pi^{+}\pi^{-}$ experiment can lead to different results, as illustrated by the triangle, diamond, and square points in Fig. 5(A).